OpenStudy (anonymous):

Solve 4x - 13 greater or equal to 20 ?

OpenStudy (anonymous):

x=13/4

OpenStudy (bahrom7893):

false

OpenStudy (nenadmatematika):

\[4x \ge33\] \[x \ge33/4\]

OpenStudy (bahrom7893):

4x-13>=20 4x>=33 x>=33/4

OpenStudy (anonymous):

\[4x - 13 \ge 20\] ??

OpenStudy (anonymous):

4x >= 20 + 13 ->x >= 33/4 = 8.25

OpenStudy (anonymous):

33/4

OpenStudy (anonymous):

sorry my mistake ! Solve \[4x - 12 \ge 20 ?\]

OpenStudy (nenadmatematika):

\[x \ge8\]